A 3.0-kg object has the following two forces acting on it: 

F1 = (16i + 12j) N 

F2 = (−10i + 22j) N 

If the object is initially at rest, determine its velocity v at t = 3.0 s. [INTEGRATE] 


1
Expert's answer
2021-11-07T19:29:53-0500

The total force is equal to the sum of the forces (bold font denotes vectors):


"\mathbf{F} = \mathbf{F_1} + \mathbf{F_2} = (16-10)\mathbf{i}+(12+22)\mathbf{j} = (6\mathbf{i}+34\mathbf{j})N"

According to the second Newton's law, the acceleration of the object is:


"\mathbf{a} = \dfrac{\mathbf{F}}{m} =\dfrac{6}{3}\mathbf{i} + \dfrac{34}{3}\mathbf{j} = \left( 2\mathbf{i}+\dfrac{34}{3}\mathbf{j}\right)m/s^2"


where "m = 3kg" is the mass of the object.

Integrating, obtain velocity:


"\mathbf{v} = \int_0^t\mathbf{a}dt = \int_0^t\left( 2\mathbf{i}+\dfrac{34}{3}\mathbf{j}\right)dt =2t \mathbf{i}+\dfrac{34t}{3}\mathbf{j}+C"

where "t" is the time and "C" is the constant of integration (velocity at "t=0" ). Since the object starts from rest, "C=0". Substituting "t = 3s", obtain:


"\mathbf{v}=2\cdot 3 \mathbf{i}+\dfrac{34\cdot 3}{3}\mathbf{j} = (6\mathbf{i}+34\mathbf{j})m/s"

Answer. "\mathbf{v}=(6\mathbf{i}+34\mathbf{j})m/s"